# Another look at inference after prediction

### v2

@@ 4.3 Improving precision of Monte Carlo simulations with control variates
The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation. Specifically, Rothery (1982) considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation. For exposition, suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$. Let $S$ denote an existing test of the same hypothesis whose power $\mu_{Z}=\mathbb{E}_{\theta}(Z)$ is known. Define $Y$ as an indicator that test $T$ rejects and $Z$ as an indicator that test $S$ rejects. For a set of $n$ simulation replicates, the baseline estimator of power for $T$ is simply $\overline{Y}=n^{-1}\sum_{i=1}^{n}Y_{i}$ while a control variate augmented estimator is given by:
@@
The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ (Rothery, 1982; Davidson and MacKinnon, 1992). The augmentation term reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$. Building on work from the survey sampling literature, a similar idea was later used for mean estimation in the semi-supervised setting by Zhang et al. (2019).
@@ 5.2 Results
Table 1: Percent biases of the various estimators. Negative values indicate underestimation, while positive values indicate overestimation of the true parameter.
Figure 2: Confidence interval length ratio of the PB inference methods compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. A ratio less than 1 indicates that the PB inference method has a confidence interval that is shorter than the confidence interval for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@ 6 Discussion
In this paper, we showed that adding a simple weight to the PPI estimator provides valid PB inference that is guaranteed to be at least as efficient as estimation using only labeled data. The weight is drawn from the work of Chen and Chen (2000) in the context of double sampling. In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that can limit its practical utility. By incorporating the proposed weight, the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference. In addition to theoretical results that support these findings, our simulation studies demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate. Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide interesting context for the PB inference problem. While we focused on least squares estimation, our results extend to other settings. The work of Chen and Chen (2000) and Angelopoulos et al. (2023a) apply to a wide variety of settings beyond linear regression, such as generalized linear models.
The work of Angelopoulos et al. (2023a) was pivotal in advancing the growing body of literature in PB inference. Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data (Zrnic and Candès, 2024b, b). Additionally, two concurrent pre-prints have noted potential losses in efficiency of PB inference approaches. The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator (Angelopoulos et al., 2023b). The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based on labeled data. In contrast, the approach of Chen and Chen (2000) minimizes the mean square error among a class of estimators with a linear augmentation term. Simultaneously, Miao et al. (2023) introduced the PoSt-Prediction Adaptive inference (PSPA) estimator that is both robust to the accuracy of the ML model and guaranteed to be at least as efficient as estimation using only labeled data. Unlike the PPI++ estimator, the PSPA estimator is designed to achieve element-wise variance reduction and was shown to outperform the PPI++ approach. The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the linear augmentation term.
Code availability
@@
Acknowledgments
The authors gratefully acknowledge funding from an NSERC Discovery Grant (RGPIN-2021-03734), the Connaught Fund, the University of Toronto Data Science Institute, and the McLaughlin Center for Genetic Research.
@@ A.1 Preliminaries
Notation. We denote the outer product of a vector $\boldsymbol{v}$ as $\boldsymbol{v}^{\otimes 2}=\boldsymbol{v}\boldsymbol{v}^{{\sf\scriptscriptstyle{T}}}$. We denote that a matrix, ${\bf{M}}$, is positive semi-definite as ${\bf{M}}\succeq{\bf{0}},$ and positive definite as ${\bf{M}}\succ{\bf{0}}.$ For any estimator $\hat{\bm{\beta}}$, we let $\text{aVar}(\hat{\bm{\beta}})$ represent its asymptotic variance-covariance matrix. We define the difference in asymptotic variance-covariance matrices between two estimators, $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{2}$, as $\Delta_{\text{aVar}}(\hat{\bm{\beta}}_{1},\hat{\bm{\beta}}_{2})=n[\text{aVar}(\hat{\bm{\beta}}_{1})-\text{aVar}(\hat{\bm{\beta}}_{2})].$
@@ A.2 Proof of Theorem 1
Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$)
@@
Proof  First, note that $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ may be decomposed as
@@
Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance-covariance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
@@
Proof  Equation (5) in the proof of Theorem 1 gives the asymptotic expansion of $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$. The standard CLT implies $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$ converges weakly to a zero-mean multivariate Gaussian distribution with variance-covariance given by
$\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$
@@
$\begin{split}n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})]=&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\left[\boldsymbol{\Sigma}_{Y,Y}+(1-\pi)\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2(1-\pi)\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.\end{split}$
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. Additionally, if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@
Proof  For scenario 1, assume that $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$. It follows that

### v3

@@ 4.3 Improving precision of Monte Carlo simulations with control variates
The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation. Specifically, Rothery [1982] considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation. For exposition, suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$. Let $S$ denote an existing test of the same hypothesis whose power $\mu_{Z}=\mathbb{E}_{\theta}(Z)$ is known. Define $Y$ as an indicator that test $T$ rejects and $Z$ as an indicator that test $S$ rejects. For a set of $n$ simulation replicates, the baseline estimator of power for $T$ is simply $\overline{Y}=n^{-1}\sum_{i=1}^{n}Y_{i}$ while a control variate augmented estimator is given by:
@@
The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ [Rothery, 1982, Davidson and MacKinnon, 1992]. The augmentation term in 3 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$. Building on work from the survey sampling literature, a similar idea was later used for mean estimation in the semi-supervised setting by Zhang et al. [2019].
@@ 5.2 Results
Table 1: Percent biases of the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), and the naive estimator ($\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$). Negative values indicate underestimation, while positive values indicate overestimation of the true parameter.
Figure 2: Confidence interval length ratio of the PB inference methods compared to baseline estimation with the labeled data ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$). A ratio less than 1 indicates that the PB inference method has a confidence interval that is shorter than the confidence interval using $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
6 UK Biobank Analysis
Central obesity, defined as excessive accumulation of body fat around abdominal region, is strongly associated with cardiometabolic diseases such as type 2 diabetes and cardiovascular disease. A major challenge in studying central obesity is the limited availability of direct body composition measurements. Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice [Zhu et al., 2022]. However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies. For example, in the UK Biobank, a prospective cohort study measuring genetic, lifestyle and health information on approximately 500,000 UK participants, DEXA-derived AFM is incompletely ascertained, particularly among populations populations historically underrepresented in genetic research [Sudlow et al., 2015]. For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
[…]
Table 2 presents point estimates, standard errors, and CI ratios (or equivalently standard errors length ratios) across the two prediction model settings for four estimators: the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), and the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$). Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators for the same parameter. For the analysis with the strong prediction model, $\widehat{Y}$ provides additional predictive power beyond the regression of $Y$ on ${\bf{X}}$, satisfying the sufficient condition in Corollary 2.1. Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In fact, all three PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$) perform similarly, yielding substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with average CI ratios below 0.85 for all methods. The analysis with the weak prediction model highlights the difference in performance between the PPI and CC methods and reinforces the findings in Corollary 2.1 and Corollary 2.2. Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with CIs 38.5% longer on average. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is no worse than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest.
Strong Prediction Model
Intercept Age Sex BMI
$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
EST -4.857 -0.003 0.813 0.173
SE 0.308 0.004 0.070 0.009
CI Ratio 1.00 1.00 1.00 1.00
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
EST -4.513 -0.007 0.904 0.164
SE 0.261 0.004 0.063 0.007
CI Ratio 0.85 0.86 0.89 0.80
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
EST -4.527 -0.007 0.900 0.165
SE 0.260 0.004 0.063 0.007
CI Ratio 0.85 0.86 0.89 0.80
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
EST -4.525 -0.006 0.905 0.163
SE 0.255 0.004 0.062 0.007
CI Ratio 0.83 0.85 0.89 0.77
Weak Prediction Model
Intercept Age Sex BMI
$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
EST -4.857 -0.003 0.813 0.173
SE 0.308 0.004 0.070 0.009
CI Ratio 1.00 1.00 1.00 1.00
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
EST -4.530 -0.006 0.835 0.165
SE 0.416 0.006 0.098 0.012
CI Ratio 1.35 1.43 1.40 1.36
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
EST -4.546 -0.006 0.834 0.165
SE 0.415 0.006 0.098 0.012
CI Ratio 1.35 1.43 1.40 1.36
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
EST -4.779 -0.004 0.819 0.171
SE 0.294 0.004 0.072 0.008
CI Ratio 0.95 0.96 1.03 0.93
Table 2: UK Biobank analysis of android fat mass based on the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$). The strong prediction model has an $R^{2}$ of 0.8, while the weak prediction model has an $R^{2}$ of 0.5. Estimates (EST), standard errors (SE), and confidence interval (CI) ratios are reported to evaluate performance. A CI ratio less than 1 indicates that a method yields a confidence interval shorter than the CI using $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@ 7 Discussion
In this paper, we showed that adding a simple weight to the PPI estimator provides valid PB inference that is guaranteed to be at least as efficient as estimation using only labeled data. The weight is drawn from the work of Chen and Chen [2000] in the context of double sampling. In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that threatens its practical utility. By incorporating the proposed weight, the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference. In addition to theoretical results that support these findings, our simulation studies and application to UK Biobank data demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate. Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide interesting context for the PB inference problem. While we focused on least squares estimation for exposition, our results extend to other settings. The work of Chen and Chen [2000] and Angelopoulos et al. [2023a] apply to a wide variety of settings beyond linear regression, including Z-estimation problems.
The work of Angelopoulos et al. [2023a] was pivotal in advancing the growing body of literature in PB inference. Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data [Zrnic and Candès, 2024b, b]. Additionally, two concurrent pre-prints have noted potential losses in efficiency of PB inference approaches. The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator [Angelopoulos et al., 2023b]. The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based on labeled data. In contrast, the approach of Chen and Chen [2000] minimizes the mean square error among a class of estimators with a linear augmentation term. Simultaneously, Miao et al. [2023] introduced the PoSt-Prediction Adaptive inference (PSPA) estimator that is both robust to the accuracy of the ML model and guaranteed to be at least as efficient as estimation using only labeled data. Unlike the PPI++ estimator, the PSPA estimator is designed to achieve element-wise variance reduction and was shown to outperform the PPI++ approach. The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the linear augmentation term.
SUPPLEMENTARY MATERIAL
Title:
The supplementary materials include proofs for theorems and additional examples for Corollary 1.1 (appendix.pdf).
Code availability:
@@
UK Biobank Dataset:
This work used data from the UK Biobank study (https://www.ukbiobank.ac.uk) and our access was approved under application 64875.
@@ A.1 Preliminaries
Notation. We denote the outer product of a vector $\bm{v}$ as $\bm{v}^{\otimes 2}=\bm{v}\bm{v}^{{\sf\scriptscriptstyle{T}}}$. We denote that a matrix, ${\bf{M}}$, is positive semi-definite as ${\bf{M}}\succeq{\bf{0}},$ and positive definite as ${\bf{M}}\succ{\bf{0}}.$ For any estimator $\hat{\bm{\beta}}$, we let $\text{aVar}(\hat{\bm{\beta}})$ represent its asymptotic variance-covariance matrix. We define the difference in asymptotic variance-covariance matrices between two estimators, $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{2}$, as $\Delta_{\text{aVar}}(\hat{\bm{\beta}}_{1},\hat{\bm{\beta}}_{2})=n[\text{aVar}(\hat{\bm{\beta}}_{1})-\text{aVar}(\hat{\bm{\beta}}_{2})].$
@@ A.2 Proof of Theorem 1
Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$).
@@
Proof.
First, note that $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ may be decomposed as
@@
Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance-covariance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
@@
Proof.
Equation (5) in the proof of Theorem 1 gives the asymptotic expansion of $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$. The standard CLT implies $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$ converges weakly to a zero-mean multivariate Gaussian distribution with variance-covariance given by
$\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}.$
@@
$\begin{split}n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})]=&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\left[\bm{\Sigma}_{Y,Y}+(1-\pi)\bm{\Sigma}_{\widehat{Y},\widehat{Y}}-2(1-\pi)\bm{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.\end{split}$
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. Additionally, if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. ∎
@@
Proof.
For scenario 1, assume that $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$. It follows that